OpenStudy (anonymous):

Explain how to solve the equation 3t + 11 = –10. Show that your solution works.

hero (hero):

3t + 11 = -10 3t = -21 t = -21/3 t = -7 3(-7) + 11 = -10 -21 + 11 = -10 -10 = -10

OpenStudy (anonymous):

I'm checkimg my answers on a quiz before I submit it soo dont hate.

hero (hero):

lol

OpenStudy (amistre64):

can i despise a little tho?

hero (hero):

I wish I had an OS when I was in school.

OpenStudy (anonymous):

no not even a little. haha

OpenStudy (amistre64):

confound it all!! lol

OpenStudy (anonymous):

ohh hush hahaha

OpenStudy (anonymous):

First you have to isolate the variable on one side of the equation (or equal sign). You can do this by subtracting 11 from both sides. This will give you the following. \[3t+11=-10\] \[3t=-10-11\] \[3t=-21\]

hero (hero):

Equation editor is just too slow for me. I've practically given up on it.

OpenStudy (anonymous):

well you just get the t to one side, then do simple math from there

OpenStudy (anonymous):

lol your name "horrible at math" doesnt give you much of a title in this section..you realize that right?? lol

OpenStudy (anonymous):

Actually, you're question was sort of common sense (things i learned in fifth grade). My username refers to the fact that I am horrible at the math i am doing now (precalculus and ap statistics) were i have A's, but not A pluses.